OpenStudy (anonymous):

im working with hyperbolas and need to know how to get the equation into x^2\a-y^2/b form. the equation im working with is 16y^2-x^2=144

ganeshie8 (ganeshie8):

you wanto change it to :- \(\large \frac{x^2}{a^2} - \frac{y^2}{b^2} = 1\)

ganeshie8 (ganeshie8):

since you're seeing \(1\) on right side, start by dividing whole thing by 144, so that u get \(1\) on right side